OpenStudy (anonymous):

pleaseee help...solve the system of equations below using elimination by addition. 5m-n=29 m-5n=-23

jhonyy9 (jhonyy9):

do you know what mean elimination by addition ,yes ?

OpenStudy (anonymous):

would it be 6

jhonyy9 (jhonyy9):

5m-n=29 m-5n=-23 ----------- elimination by addition mean that you multiplie first or second equation by a number so that if you add term by term one from these will be eliminated for example x +3y = 2 2x -y =1 / *3 -------------- for in this case will be eliminated one from x and y so you can multiplie second equation by 3 and will get 6x -3y =3 than will be x +3y =2 6x -3y =3 ------------- adding term by term 7x 0 =5 7x=5 x=5/7 do you understand it now ?

jhonyy9 (jhonyy9):

so this was just an example

jhonyy9 (jhonyy9):

do you understand it now sure right ? than you can tell me what you can doing with your exercise ?

jhonyy9 (jhonyy9):

what equation you need multiplie by how many for will be eliminated one from these terms

jhonyy9 (jhonyy9):

5m-n=29 m-5n=-23 so for being eliminated m you need multiplie second equation by how many ?

jhonyy9 (jhonyy9):

so you dont like cooperate anything ?

OpenStudy (anonymous):

so 5m-5n=52

jhonyy9 (jhonyy9):

no i think that you not have understood it nothing from my above example

OpenStudy (anonymous):

its loooking for a pair

jhonyy9 (jhonyy9):

5m -n =29 m -5n = -23 so now if you check it for eliminate the m adding you need multiplie second equation by -5 so than what will get ?

jhonyy9 (jhonyy9):

please try it alone

jhonyy9 (jhonyy9):

5m-n=29 m -5n = -23 *(-5) ---

OpenStudy (anonymous):

m-5n=-115

jhonyy9 (jhonyy9):

will get -5m +25n = 115

jhonyy9 (jhonyy9):

how you have got this ?

jhonyy9 (jhonyy9):

you need multiplie the both sides

jhonyy9 (jhonyy9):

ok ?

jhonyy9 (jhonyy9):

so than will be 5m -n =29 -5m +25n =115 ---------------- so than now adding term by term will get ?

OpenStudy (anonymous):

20n=115

OpenStudy (anonymous):

24n=144

OpenStudy (anonymous):

answer is 6

jhonyy9 (jhonyy9):

adding will be eliminated the m -n+25n=29 +115 yes now is right

jhonyy9 (jhonyy9):

ok than you know that n=6 you need getting again m

jhonyy9 (jhonyy9):

5m-n=29 5m-6=29 continue ...

OpenStudy (anonymous):

10m-7n=58

jhonyy9 (jhonyy9):

no there is 5m - 6=29 calculi from this the m

jhonyy9 (jhonyy9):

add to both sides 6 5m-6+6=29+6 5m=35 m=35/5 m=7

jhonyy9 (jhonyy9):

ok ?

jhonyy9 (jhonyy9):

hope so much that you understand it sure easy good luck was my pleasure bye

OpenStudy (anonymous):

so the points would be 6,7 right or 7,6 thank you so much